# Large Language Models Engineer Too Many Simple Features for Tabular Data

### v1

@@ 6 Conclusion
In conclusion, the contributions of our work are a method to detect bias in LLMs and evidence that a bias toward simple operators exists. Future work should explore methods to circumvent this bias. Promising methods are in-context learning (e.g., prompt tuning) or fine-tuning the LLM to favor optimal operators. In the long term, after identifying and addressing biases in LLMs, we can fully liberate ourselves from manual feature engineering. This will allow us to leverage LLMs as reliable and efficient automated data science agents for tabular data.
@@
Our study on the bias of LLMs still has limitations because it is the first of its kind for automated data science. We detect a bias but do not support practitioners to determine why an LLM might be biased. Similarly, given how frontier LLMs are trained and deployed, we focused on assessing bias in the models’ outputs rather than internal mechanisms. Lastly, our study is limited to four LLMs, while practitioners can also choose from many other potentially biased LLMs. Our findings and proposed method do not create any negative societal impacts. Instead, both can have positive societal impacts because they increase our understanding and (mis)trust when using large language models for feature engineering.
@@
Acknowledgments
We acknowledge funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under SFB 1597 (SmallData), grant number 499552394. This research was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under grant number 417962828. Frank Hutter acknowledges the financial support of the Hector Foundation.
@@ A.1 Prompt Templates
Figure 7: Our System Prompt. The contents of the full system prompt, which is send to the LLM before the instructions to generate new features for a given dataset.
Figure 8: Our Instruction Prompt. The contents of the instruction prompt on the example of the blood-transfusion-service-center dataset. This prompt is send every time the LLM is instructed to generate a feature for a given dataset. The order of operators is shuffled according to the explanations in Section 4.
A.2 Feedback Loop
@@ Appendix B List of Operator and their Categories
Table 4: Simple Operators.
@@
Table 5: Complex Operators.
@@ Appendix C Datasets
Table 6: Benchmark Datasets The table contains all datasets from the AutoML benchmark (Gijsbers et al., 2024) that we used in our experiments. We selected dataset based on our constraints defined in Section 4. All datasets listed in this Table were tested for memorization of the LLMs.
@@
Appendix D Large Language Models
Table 7: Model API Versions. The full model versions as specified by the respective API provider.
Model Model Version
gpt-4o-mini gpt-4o-mini-2024-07-18
gemini-1.5-flash gemini-1.5-flash-001
mistral7b-v0.3 Mistral-7B-Instruct-v0.3 (FP16 Quantization)
Appendix E Memorization Test Results
Table 8: Memorization Tests Results GPT and Llama. We present the results of all three memorization tests (Bordt et al., 2024), (1) row completion test (r.c.), (2) feature completion test (f.c.) and first token test (f.t.) for gpt-4o-mini and llama3.1-8b. For each dataset and each test, the number of runs which imply signs of memorization are listed. Each test ran 25 tries per dataset. Sometimes, the LLM failed to match the expected outcome sequences required by the tests (noted as -). If a prior language model exhibited signs of memorization for a dataset, the tests were not further conducted for subsequent models (noted as X)
@@ Appendix E Memorization Test Results
Table 9: Memorization Tests Results Mistrial and Gemini. We present the results of all three memorization tests (Bordt et al., 2024), (1) row completion test (r.c.), (2) feature completion test (f.c.) and first token test (f.t.) for mistral7b-v0.3 and gemini-1.5-flash. For each dataset and each test the number of runs which imply signs of memorization are listed. Each test ran 25 tries per dataset. Sometimes, the LLM failed to match the expected outcome sequences required by the tests (noted as -). If a prior language model exhibited signs of memorization for a dataset, the tests were not further conducted for subsequent models (noted as X)
@@
Appendix F Predictive Accuracy Results
Table 10: Predictive Performance of Feature Engineering. We show the average and standard deviation of the ROC AUC scores for all folds for all datasets. Base represents the baseline score without feature engineering. The scores for the four large language models additionally contain the average over all 5 shuffles of operator order in the instructions prompt. For each fold the respective method generated 20 new features. Features were only added to the method if it improved the feedback scores described in A.

### v2

@@ 6. Conclusion
In conclusion, the contributions of our work are a method to detect bias in LLMs and evidence that a bias toward simple operators exists. The findings of this method underscore the necessity to further strengthen LLMs to truly unlock their potential for tabular data problems. Our work highlights the importance of developing methods to mitigate bias in downstream applications. Promising methods for future work to explore are in-context learning (e.g., prompt tuning) or fine-tuning the LLM to favor optimal operators. In the long term, after identifying and addressing biases in LLMs, we can fully liberate ourselves from manual feature engineering. This will allow us to leverage LLMs as reliable and efficient automated data science agents for tabular data.
@@
Our study on the bias of LLMs still has limitations because it is the first of its kind for automated data science. We detect a bias but do not support practitioners in determining why an LLM might be biased. Similarly, given how frontier LLMs are trained and deployed, we focused on assessing bias in the models’ outputs rather than internal mechanisms. Lastly, our study is limited to four LLMs, while practitioners can also choose from many other potentially biased LLMs. Our findings and proposed method do not create any negative societal impacts. Instead, both can have positive societal impacts because they increase our understanding and (mis)trust when using large language models for feature engineering.
@@ A.1. Prompt Templates
A.2. Feedback Loop
@@ Appendix B List of Operator and their Categories
Table 4. Simple Operators.
@@
Table 5. Complex Operators.
@@ Appendix C Datasets
Appendix D Large Language Models
Table 6. Model API Versions. The full model versions as specified by the respective API provider.
Model Model Version
gpt-4o-mini gpt-4o-mini-2024-07-18
gemini-1.5-flash gemini-1.5-flash-001
mistral7b-v0.3 Mistral-7B-Instruct-v0.3 (FP16 Quantization)
Appendix E Memorization Test Results
Appendix F Predictive Accuracy Results
Appendix G Additional Experiments
To further solidify the results of our study we conducted some additional experiments.
G.1. Statistical Significance
We test the statistical significance of our results on predictive accuracy across all benchmark datasets and methods. The results of these tests are presented in the critical difference diagram in Figure 8. As visible in this diagram, OpenFE has the highest rank across all methods, outperforming all LLM-based methods. GPT-4o-mini and Gemini-1.5-flash exhibit no statistical difference and are last in ranks, matching our findings from Table 3 and Figure 5. Further, the similarity in performance between Llama3.1-8b and Mistral-7b-v0.3 is further solidified.
Figure 8. Critical Difference Plots for Test Scores. Mean rank of the methods (lower is better). Methods connected by a bar are not significantly different.
G.2. Feature Selection Frequencies
We evaluate the frequencies of selected features per dataset. For each dataset we calculate the frequencies with which each feature is selected over all feature generation steps. As presented in Figure 11 the LLM is relatively certain which features to select, indicated by high frequencies for few features on most datasets.
Figure 9. Our System Prompt. The contents of the full system prompt, which is sent to the LLM before the instructions to generate new features for a given dataset.
Figure 10. Our Instruction Prompt. The contents of the instruction prompt on the example of the blood-transfusion-service-center dataset. This prompt is send every time the LLM is instructed to generate a feature for a given dataset. The order of operators is shuffled according to the explanations in Section 4.
Table 7. Benchmark Datasets The table contains all datasets from the AutoML benchmark (Gijsbers et al., 2024) that we used in our experiments. We selected dataset based on our constraints defined in Section 4. All datasets listed in this Table were tested for memorization of the LLMs.
@@ G.2. Feature Selection Frequencies
Table 8. Memorization Tests Results GPT and Llama. We present the results of all three memorization tests (Bordt et al., 2024), (1) row completion test (r.c.), (2) feature completion test (f.c.) and first token test (f.t.) for gpt-4o-mini and llama3.1-8b. For each dataset and each test, the number of runs which imply signs of memorization are listed. Each test ran 25 tries per dataset. Sometimes, the LLM failed to match the expected outcome sequences required by the tests (noted as -). If a prior language model exhibited signs of memorization for a dataset, the tests were not further conducted for subsequent models (noted as X)
@@
Table 9. Memorization Tests Results Mistrial and Gemini. We present the results of all three memorization tests (Bordt et al., 2024), (1) row completion test (r.c.), (2) feature completion test (f.c.) and first token test (f.t.) for mistral7b-v0.3 and gemini-1.5-flash. For each dataset and each test the number of runs which imply signs of memorization are listed. Each test ran 25 tries per dataset. Sometimes, the LLM failed to match the expected outcome sequences required by the tests (noted as -). If a prior language model exhibited signs of memorization for a dataset, the tests were not further conducted for subsequent models (noted as X)
@@
Figure 11. Feature Selection Frequency per Dataset. We present the frequencies with which each feature in a dataset is selected by the LLM over all feature generation steps. In most cases the LLM selects few features with a high frequency repeatedly.
Table 10. Predictive Performance of Feature Engineering. We show the average and standard deviation of the ROC AUC scores for all folds for all datasets. Base represents the baseline score without feature engineering. The scores for the four large language models additionally contain the average over all 5 shuffles of operator order in the instructions prompt. For each fold the respective method generated 20 new features. Features were only added to the method if it improved the feedback scores described in A.